OpenStudy (anonymous):

7y - 3 + 3y + 3 = 90. Please Show Steps If you don't care 7y - 3 + 3y + 3 = 90. Please Show Steps If you don't care @Mathematics

OpenStudy (agreene):

7y-3+3y+3 = 90 start by combining likes on the left 10y=90 divide by 10 on both sides y=90/10 y=9

OpenStudy (saifoo.khan):

7 + 3 =11 wow, hats off to me! :D

OpenStudy (anonymous):

Thanks man

OpenStudy (agreene):

i was just typing something along those likes saifoo